A current of 10 Ampere exist in a line of 0.15 ohms resistance compute the rate of production of thermal energy in watts
1
Expert's answer
2016-11-14T10:05:07-0500
The electric power can be calculated as:
P=I^2*R=(10A)^2*0.15 Ω=15 W